Need an equasion and answer for:
The length of a rectangle is 6 meters more than 3 x the width the perimeter is 44 meters find the length and width.

Answer by pie90010(44)

so you have a lot to work with here
P = 44
Perimeter in general equals lenght * width
L = 6 + 3W
W = W
so just make 1 big equation by adding all 4 sides
8W + 12 = 44
8W = 32
W = 4
now that you have the width
find the length by plugging it in to the original L equation
L = 6 + 3w
L = 6 + 3(4)
L = 6 + 12
L = 18
So the answer would be
L = 18
W = 4